OpenStudy (anonymous):

write the expression as a complex number in standard form. (2+4i)+(6-i)

OpenStudy (mathstudent55):

Here, the operation to the left of the second set of parentheses is plus. There is no operation to the left of the first set of parentheses. This means, step 1 is drop both sets of parentheses. Step 2 is add like terms.

OpenStudy (anonymous):

so it would be 8-3i

OpenStudy (mathstudent55):

4i - i is????

OpenStudy (mathstudent55):

2 + 6 = 8, you're correct there

OpenStudy (anonymous):

because its a negative 3i i subtracted it

OpenStudy (mathstudent55):

4i is positive -i is negative 4i - i is similar to 4 - 1. What is 4 - 1?

OpenStudy (anonymous):

3

OpenStudy (mathstudent55):

Right, so just like 4 - 3 = 1, so 4i - i = 3i Remember that i = 1i, so 4i - i = 4i - 1i = 3i

OpenStudy (anonymous):

so it would be 8+3i

OpenStudy (espex):

Yes.

OpenStudy (anonymous):

thank you can you help me with another problem